OpenStudy (anonymous):

Simplify 27x^3/9x^2. Could someone explain it to me because it say on my worksheet that it's 3x simplified but I keep getting 9x.

OpenStudy (anonymous):

Hey first of all: \[\huge \color{blue}{\textbf{Welcome To Openstudy...}}\]

OpenStudy (anonymous):

See: For x part use this: \[\large \color{green}{\frac{x^a}{x^b} = x^{a - b}}\]

OpenStudy (anonymous):

Here as you are seeing a = 3 and b = 2, so proceed further and tell me what are you getting??

OpenStudy (anonymous):

I don't understand. The way my teacher showed it we had to put it in standard form and then cross out what both sides have.

OpenStudy (anonymous):

So, let us do it by one another method: \[\large \color{blue}{x^3 = x \times x \times x}\] And: \[\large \color{red}{x^2 = x \times x}\] Am I right??

OpenStudy (anonymous):

See, if you have any difficulty then ask me directly and remember I am not your teacher but a Friend, so let everything clear here only..

OpenStudy (anonymous):

It's more like:  And alright, thank you.

OpenStudy (anonymous):

And I am seeing that you got "x" in your answer that is right.. But you got "9" with it on which we have to work more.. Waiting for your reply..

OpenStudy (anonymous):

Yes, I cross out all the threes because they are the same on both sides but I don't know if I should simplify the 9. I got the X's correctly.

OpenStudy (anonymous):



OpenStudy (anonymous):

3x?

OpenStudy (anonymous):

Yep.. Not getting this part??

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

Here can you tell what are getting cancelled??

OpenStudy (anonymous):

The three's and two x's on both sides.

OpenStudy (anonymous):

\[9x\] That.

OpenStudy (anonymous):

Hey wait: \[\huge \frac{3 \times 9 \times x \times x \times x}{9 \times x \times x}\] This is your original question I think..

OpenStudy (anonymous):

Am I right or wrong??

OpenStudy (anonymous):

Well originally it was: 3 * 9 * x * x * x 3 * 3 * x * x I don't think that I can just multiply the three's on the lower part. I would have to cancel them out.

OpenStudy (anonymous):

But I think you might be right. That I multiply the bottom three's and cancel out the nines.

OpenStudy (anonymous):

See, on lower side why are you writing 9 as 3*3??

OpenStudy (anonymous):

That's how it was shown on my paper how to do it. It didn't give me any other explanation.

OpenStudy (anonymous):

Keep that 9 as it will be simple for you to understand.. See, on upper side you can write 27 as : 3*9 And keep 9 itself on lower side.. So you will get: \[\large \frac{3 \times 9 \times x \times x \times x}{9 \times x \times x}\] Now cancel whatever you can cancel..

OpenStudy (anonymous):

So just cancel out the nines and two x's on both sides.

OpenStudy (anonymous):

Yep so are you left with: \[\large 3x\]

OpenStudy (anonymous):

I get it now. Thanks so much.